# Markov Decision Processes                                  Under External Temporal Processes

### v2

@@ Appendix A Proofs of lemmas
Lemma B.1 (Generalization Lemma for Markov chains, Lemma 24 of Lazaric et al. (2012)).
Let $\tilde{\mathcal{F}}$ be the $d$-dimensional class of linear functions truncated at threshold $B$. Let $\left(X_{t}\right)_{t\in[n]}$ be a sequence of samples from an ergodic, aperiodic exponentially $\beta$-mixing Markov chain with parameters $\bar{\beta},b,\kappa$, arbitrary initial distribution, and stationary distribution $\rho$. If the first $\tilde{n}$ samples are discarded and only the last $n-\tilde{n}$ samples are used, then with probability at least $1-\delta$, the empirical and $l_{2}(\rho)$ norm of the any function $\tilde{f}\in\tilde{\mathcal{F}}$ are related as,
$||\tilde{f}||-2||\tilde{f}||_{X_{1:n}}$ $\leq\epsilon(\delta),$
$||\tilde{f}||_{X_{1:n}}-2\sqrt{2}||\tilde{f}||$ $\leq\epsilon(\delta),$
$\text{for }\epsilon(\delta)=12B\sqrt{\frac{2\Lambda(n-\tilde{n},d,\delta)}{n-\tilde{n}}\max\left\{\frac{\Lambda(n-\tilde{n},d,\delta)}{1},1\right\}^{\frac{1}{\kappa}}},$
@@
Lemma B.2 (Lemma 4 of Lazaric et al. (2012)).
@@
comes from a simpler version of Lemma B.1 without the additional initial $\tilde{n}$ samples for burn-in of the Markov chain.
The generalization Lemma B.1, combined with the lower bound on the eigenvalues and our results in Section 6, give rise to the proof of the following result on the sample complexity of policy evaluation, detailed in Section 7.
B.1. Theorem 7.1
Assume that the policy $\pi$ induces from the MDP $\mathcal{M}_{X}$ a $\beta$-mixing Markov chain with parameters $\beta$, $b$, $\kappa$. Let $\bar{s}_{1},\dots,\bar{s}_{N+\tilde{N}}$ be a sample path obtained using the policy $\pi$. Suppose the first $\tilde{N}$ samples are discarded for Markov chain burn-in, and the remaining $N$ are used to compute the truncated least-squares path-wise estimate $\tilde{V}$ of the true value function $V$. Then, provided the number of discarded samples is $\tilde{N}=\left(\frac{1}{b}\log\frac{2e\bar{\beta}n}{\delta}\right)^{\frac{1}{\kappa}}$, with probability at least $1-\delta$,
$||\tilde{V}-V||_{\rho}$ $\leq\frac{4\sqrt{2}}{\sqrt{1-\gamma^{2}}}\left(\frac{3}{(1-\gamma^{2})}\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)+||\Pi^{trunc}V-\Pi V||_{\rho}\right)$
@@
with $\Lambda_{1}(N,d,\delta)$ and $\Lambda_{2}(N,\delta)$ as defined in Lemma B.1, and $\alpha^{*}$ such that $\Pi V=\sum_{i}\alpha^{*}_{i}\varphi_{i}$.

### v3

@@ Appendix B Proofs of lemmas
Lemma C.1 (Generalization Lemma for Markov chains, Lemma 24 of Lazaric et al. (2012)).
Let $\widetilde{\mathcal{F}}$ be a $d$-dimensional class of linear functions truncated at threshold $B$. Let $\left(X_{t}\right)_{t\in[n]}$ be a sequence of samples from an ergodic, aperiodic exponentially $\beta$-mixing Markov chain with parameters $\bar{\beta},b,\kappa$, arbitrary initial distribution, and stationary distribution $\rho$. If the first $\widetilde{n}$ samples are discarded and only the last $n-\widetilde{n}$ samples are used, then with probability at least $1-\delta$, the empirical and $l_{2}(\rho)$ norm of the any function $\widetilde{f}\in\widetilde{\mathcal{F}}$ are related as,
$\|\widetilde{f}\|-2\|\widetilde{f}\|_{X_{1:n}}$ $\leq\epsilon(\delta),$
$\|\widetilde{f}\|_{X_{1:n}}-2\sqrt{2}\|\widetilde{f}\|$ $\leq\epsilon(\delta),$
$\text{for }\epsilon(\delta)=12B\sqrt{\frac{2\Lambda(n-\widetilde{n},d,\delta)}{n-\widetilde{n}}\max\left\{\frac{\Lambda(n-\widetilde{n},d,\delta)}{1},1\right\}^{\frac{1}{\kappa}}},$
@@
Lemma C.2 (Lemma 4 in Lazaric et al. (2012)).
@@
comes from a simpler version of Lemma C.1 without the additional initial $\widetilde{n}$ samples for burn-in of the Markov chain.
The generalization Lemma C.1, combined with the lower bound on the eigenvalues and our results in Section 5, give rise to the proof of the following result on the sample complexity of policy evaluation, detailed in Section 6.
C.1. Theorem 6.1
Assume that the policy $\pi$ induces from the MDP $\mathcal{M}_{X}$ a $\beta$-mixing Markov chain with parameters $\beta$, $b$, $\kappa$. Let $\bar{s}_{1},\dots,\bar{s}_{N+\widetilde{N}}$ be a sample path obtained using the policy $\pi$. Suppose the first $\widetilde{N}$ samples are discarded for Markov chain burn-in, and the remaining $N$ are used to compute the truncated least-squares path-wise estimate $\widetilde{V}$ of the true value function $V$. Then, provided the number of discarded samples is $\widetilde{N}=\left(\frac{1}{b}\log\frac{2e\bar{\beta}n}{\delta}\right)^{\frac{1}{\kappa}}$, with probability at least $1-\delta$,
$\left\|\widetilde{V}-V\right\|_{\rho}$ $\leq\frac{4\sqrt{2}}{\sqrt{1-\gamma^{2}}}\left(\frac{3}{(1-\gamma^{2})}\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)+\left\|\Pi^{trunc}V-\Pi V\right\|_{\rho}\right)$
@@
with $\Lambda_{1}(N,d,\delta)$ and $\Lambda_{2}(N,\delta)$ as defined in Lemma C.1, and $\alpha^{*}$ such that $\Pi V=\sum_{i}\alpha^{*}_{i}\varphi_{i}$.